OpenStudy (anonymous):

ok i got this far but i dont what to do next.. : 3d + 4 = 2 + 3d - 1/2 3d +2 = + 3d - 0.5 but for some reason i got 2.5 HELPPPP PLEASE

OpenStudy (anonymous):

what was the original question?

OpenStudy (anonymous):

3d + 4 = 2 + 3d - 1/2

OpenStudy (anonymous):

based on what you posted solving for d we get 2 + 0.5= 3d - 3d 2.5 = 0 which obviously is not correct

OpenStudy (anonymous):

The question wants you to solve for d correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then I can rearrange the equation to 3d + 2.5 = 3d This question appears to have no solution from my perspective which is once again proven through solving for d and getting 2.5 = 3d - 3d 2.5 = 0 which is not true. I am sorry but unless you either have more information to give or a correction to make I personally cannot help you anymore my answer would be that d does not exist

OpenStudy (anonymous):

thank you